1. According to the Appendix at the back of the text:
ΔH∘f(kJ/mol) S∘(J/mol•K)
C3H8(g) –103.8 269.9
CO2(g) –393.5 213.8
H2O(l) –285.8 69.9
O2(g) 0.0 205.0
Using the standard heats of formation and the absolute entropies, calculate ΔG∘ (in units of kJ) for the following reaction at 300 K. answer in KJ.
C3H8(g) + 5O2(g)=3CO2(g)+4H2O(l)
1
Expert's answer
2020-02-26T05:02:06-0500

Hr = 3*(-393.5)+4*(-285.8)-1*(-103.8) = -2221 kJ

Sr = 3*213.8+4*69.9-1*269.9-5*205 = -373 kJ

G = H - T*S = - 2221 - 300*(-0.373) = - 2109.1 kJ